You have a RC circuit which contains a capacitor of 0.02 F and a resistor of 350 Ohm. Calculate how 

long it would take (in seconds) for the capacitor to be charged to 89% of its full voltage.


1
Expert's answer
2022-06-10T02:31:02-0400
"Q=Q_{\max}(1-e^{-t/RC})"

"1-e^{-t\/RC}=0.89"

"e^{-t\/RC}=0.11"

"t=-RC\ln 0.11\\
=-350*0.02*\ln 0.11=15\:\rm s"